A number is as much less than 75 as it is greater by 75. Find the number
Answer by ankor@dixie-net.com(22740)

A number is as much less than 75 as it is greater by 75. Find the number
75 - n = n - 75
75 + 75 = n + n
150 = 2n
n = 75